# Conformal Prediction for Ensembles: Improving Efficiency via Score-Based Aggregation

### v2

@@ 4.2 UCI Regression Tasks
Table 2: Average coverages across tasks for $\alpha=0.05$ are shown in the top rows and average prediction set lengths in the bottom rows, both were assessed over a batch of i.i.d. test samples (10% of the datasets). Standard deviations and means were computed across 10 randomized draws of the training, calibration, and test sets. Note that, while the single-stage prediction regions are the smallest, they fail to achieve the desired coverage level and are, thus, precluded from comparison.
Dataset Methods Linear Model LASSO Random Forest Neural Net $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA (Single-Stage) CSA
3droad Coverage 0.947 (0.001) 0.947 (0.001) 0.948 (0.001) 0.95 (0.0) 0.945 (0.001) 0.915 (0.0) 0.948 (0.0) 0.926 (0.001) 0.95 (0.0)
Length 71.583 (0.325) 71.576 (0.331) 5.362 (0.006) 56.298 (0.602) 53.166 (0.482) 29.04 (0.245) 51.158 (0.015) 1.175 (0.036) 1.361 (0.031)
parkinsons Coverage 0.948 (0.002) 0.951 (0.001) 0.935 (0.005) 0.942 (0.001) 0.946 (0.002) 0.878 (0.006) 0.937 (0.005) 0.857 (0.019) 0.961 (0.004)
Length 37.058 (0.599) 37.349 (0.34) 0.876 (0.11) 18.912 (0.931) 17.918 (0.814) 8.738 (0.427) 23.945 (0.039) 0.465 (0.196) 0.588 (0.188)
pol Coverage 0.942 (0.007) 0.944 (0.007) 0.951 (0.002) 0.952 (0.001) 0.936 (0.006) 0.9 (0.009) 0.949 (0.003) 0.889 (0.007) 0.95 (0.003)
Length 97.558 (0.808) 97.476 (1.0) 19.437 (0.374) 41.775 (1.306) 38.033 (1.092) 27.596 (0.77) 63.011 (0.894) 1.357 (0.019) 1.654 (0.032)
pumadyn32nm Coverage 0.932 (0.008) 0.933 (0.008) 0.943 (0.004) 0.924 (0.011) 0.917 (0.008) 0.886 (0.013) 0.935 (0.008) 0.863 (0.018) 0.926 (0.006)
Length 3.865 (0.043) 3.827 (0.061) 1.158 (0.005) 1.145 (0.012) 1.239 (0.011) 1.089 (0.013) 2.54 (0.029) 2.355 (0.124) 2.767 (0.076)
slice Coverage 0.951 (0.001) 0.948 (0.001) 0.948 (0.0) 0.953 (0.003) 0.947 (0.0) 0.912 (0.0) 0.951 (0.002) 0.91 (0.004) 0.95 (0.001)
Length 35.242 (0.099) 35.863 (0.057) 2.749 (0.059) 8.596 (0.412) 8.315 (0.382) 5.378 (0.216) 20.401 (0.162) 1.114 (0.025) 1.388 (0.02)
@@ 4.3 CSA Predict-Then-Optimize
We now study the use of CSA for a real-world predict-then-optimize traffic routing task, as studied in the conformal predict-then-optimize paper [36]. In this task, a time series of $T$ previously observed precipitation readings are used to predict future precipitations and, in turn, anticipated traffic, as fully described in Appendix E. We consider the traffic routing problem for a fixed source-target pair $(s,t)$ over the graph of Manhattan, where $|\mathcal{V}|=4584$ and $|\mathcal{E}|=9867$. Formally, such a problem seeks
$w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}\widehat{c}^{T}w$ (5)
[…]
@@
We then consider two probabilistic models for traffic prediction, namely one based on the classical probabilistic Lagrangian integro-difference approach (STEPS) of [46] and one on the modern latent diffusion model (LDM) approach of [47]. As a result of the higher inference cost of the latter, we consider the setup where $J_{1}>J_{2}$, specifically with $J_{1}=4$ and $J_{2}=1$, which additionally serves to empirically highlight the flexibility of non-uniform sampling from predictors discussed in Section 3.2. As discussed in Section 4, the alternate aggregation strategies do not lend themselves for use in this setting. We, therefore, only compare CSA to the best model selection strategy, specifically by comparing to the separate conformalizations of the two constituent predictors. Each model was conformalized with Equation 3 taken to be the score, where $d_{k}$ was set to be the standard 2-norm across both predictors, such that $s_{1}(x,c)=\min_{j\in 1,...,J_{1}}||\widehat{c}_{1j}-c||_{2}$ for $q_{1}(C\mid x)$ and similarly for $q_{2}(C\mid x)$. We here evaluate the methods using the expected suboptimality gap proportion, $\Delta_{\%}=\mathbb{E}_{X}[\Delta(X,C(X))/\min_{w}f(w,C(X))],$ where $\Delta$ is defined as discussed in Section 2.3. This quantity measures the conservatism of the produced robust optimal value and is bounded in $[0,1]$ in this setting.
Experiments were conducted with $|\mathcal{D}_{C}|=200$. The suboptimality was then computed across 100 i.i.d. test samples. To, therefore, assess the improvement offered by CSA, we conduct two paired t-tests, with hypotheses
$H_{0}:\Delta^{(\mathrm{CSA})}_{\%}=\Delta^{(\mathrm{STEPS})}_{\%}\quad H_{1}:\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$
and similarly for $\Delta^{(\mathrm{CSA})}_{\%}$ and $\Delta^{(\mathrm{LDM})}_{\%}$. The coverages and results of such t-tests are provided in Table 3, from which we find that CSA significantly reduces the suboptimality after accounting for Bonferroni multiple testing. We see that, while conformalization of either of the two views individually already produces the desired coverage, CSA produces more informative prediction regions, and hence less conservative robust upper bounds.
Table 3: Coverages across tasks for $\alpha=0.05$ are shown in the left table across both the individually conformalized and CSA fused approaches. P-values of the paired t-tests comparing gap proportions ($\Delta_{\%}$) are shown in the right table. Both were computed over 100 i.i.d. test samples.
STEPS LDM CSA
0.981 0.962 0.968
$H_{1}$ p-value
$\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$ $3.61\times 10^{-4}$
$\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{LDM})}_{\%}$ $9.50\times 10^{-4}$
@@ 5 Discussion
We have presented CSA, a framework for producing informative prediction regions in ensemble predictor pipelines. We additionally demonstrated the generality of the proposed pipeline in settings of both classification and robust predict-then-optimize regression tasks. This work suggests many directions for extension. One point of interest is the integration of CSA with other downstream regression applications. For instance, [48] proposed an end-to-end differentiable extension to [36]: extending CSA for integration to their framework would be of great interest. Additionally, given the prevalence of late-stage fusion in robotics applications, an exciting applied avenue of research would investigate the use of CSA in settings of robust control.
@@ Appendix C CSA Predict-Then-Optimize Visual Walkthrough
Notice that, for any fixed $\vec{j}$, this is a standard convex optimization problem with a convex feasible region. We illustrate how the feasible region would be constructed for a fixed $\vec{j}$ in Figure 8. Note that the construction of this feasible is never explicitly done in practice and is only implicitly used by convex solver routines in practice.
@@
We can, therefore, then solve Equation 12 over all possible $\vec{j}\in\mathcal{J}$ and aggregate the maxima to compute $c^{*}$.
Appendix D Additional UCI Results
We consider those regression tasks from the UCI repository [45] that have at least 1,000 samples. The complete list of tasks considered herein is provided in Table 4.
Table 4: Tasks considered for empirical validation of prediction region efficiency and coverage, along with their relevant dimensions. All tasks are scalar regression tasks, i.e. $y\in\mathbb{R}$.
Dataset Observations Input Dimension
3droad 434874 3
airfoil 1503 5
bike 17379 17
parkinsons 5875 20
pol 15000 26
pumadyn32nm 8192 32
slice 53500 385
The complete collection of results is presented in Table 5. As in the main text, we see reduced conservatism of the CSA prediction regions compared to alternatives. We additionally notice that the single-stage CSA approach fails to retain the coverage guarantees.

### v3

@@ 4.1 Classification Tasks
Table 1: Classification results are shown across tasks for $\alpha=0.10$, $\alpha=0.05$, and $\alpha=0.01$, with coverages in the top (grey) and average prediction set sizes (white) in the bottom of each row. Both were assessed over a batch of i.i.d. test samples (15% of the validation set from ImageNet). Standard deviations and means were computed across 10 randomized draws of the calibration and test sets.
Dataset/$\boldsymbol{\alpha}$ ResNet VGG DenseNet VFCP $\boldsymbol{\mathcal{C}^{M}}$ $\boldsymbol{\mathcal{C}^{R}}$ $\boldsymbol{\mathcal{C}^{U}}$ Ensemble CSA
ImageNet 0.901 (0.005) 0.902 (0.003) 0.902 (0.003) 0.899 (0.004) 0.938 (0.003) 0.909 (0.004) 0.9 (0.004) 0.899 (0.004) 0.9 (0.003)
$(\alpha=0.10)$ 137.004 (1.98) 136.116 (2.206) 120.096 (2.427) 46.063 (1.089) 87.337 (1.604) 82.746 (1.692) 131.856 (2.378) 69.123 (1.317) 34.006 (0.924)
$(\alpha=0.05)$ 0.95 (0.003) 0.949 (0.004) 0.952 (0.002) 0.95 (0.003) 0.975 (0.002) 0.954 (0.004) 0.95 (0.003) 0.949 (0.002) 0.95 (0.003)
220.022 (2.072) 229.523 (3.076) 208.658 (2.016) 78.108 (2.004) 166.933 (2.157) 143.323 (2.932) 220.491 (2.773) 112.161 (2.115) 59.574 (3.382)
$(\alpha=0.01)$ 0.99 (0.001) 0.991 (0.001) 0.989 (0.002) 0.99 (0.001) 0.997 (0.001) 0.991 (0.002) 0.99 (0.002) 0.99 (0.002) 0.99 (0.002)
491.952 (6.353) 726.028 (12.157) 459.399 (6.739) 194.691 (4.579) 580.592 (7.715) 532.155 (24.829) 559.188 (7.07) 299.453 (6.526) 201.32 (46.509)
We first study the predictive efficiency of the aforementioned methods on the ImageNet classification task [43]. In particular, an ensemble was constructed from three separately trained deep learning architectures, namely ResNet-50, VGG-11, and DenseNet-121. Conformalization on the individual models was performed using the standard classification score function across all approaches, namely $s(x,y)=\sum_{j=1}^{l}\widehat{f}(x)_{\pi_{j}(x)}$ where $y=\pi_{l}(x)$ and $\pi(x)$ is the permutation of $\{1,\ldots,|\mathcal{Y}|\}$ that sorts $\widehat{f}\left(x\right)$ from most to least likely. Here, the “Ensemble” score was computed with the same $s(x,y)$, replacing $f_{k}(x)$ with the ensemble average probability, i.e. $\mu(x)_{j}:=\sum_{k}(f_{k}(x))_{j}/K$. Calibration was performed using 85% of the ImageNet test set and assessment of the coverage and interval lengths on the remaining 15%, with 10 trials conducted over randomized draws of these calibration and test sets. A $25/75\%$ split was used for $\mathcal{D_{C}}^{(1)}$-$\mathcal{D_{C}}^{(2)}$. The results are presented in Table 1; the full results across additional $\alpha$ is given in Appendix F. We see that all the approaches exhibit the desired coverages across $\alpha$. However, CSA consistently produces significantly smaller prediction regions than both the individually conformalized models and alternate aggregation strategies.
4.2 Regression Tasks
We now similarly study the predictive efficiency of CSA across a suite of regression tasks from [42]. The data for each task were split with 50/45/5% for training, calibration, and testing for coverage and interval lengths, with five trials conducted over randomized selections of such sets. A $5/95\%$ split was used for $\mathcal{D_{C}}^{(1)}$-$\mathcal{D_{C}}^{(2)}$. The problem setup was replicated from [29], in which four prediction methods were ensembled, namely an OLS model, a LASSO linear model, a random forest (RF), and an XGBoost model. A residual function was used as the score across all methods, namely $s(x,y)=|\widehat{f}(x)-y|$. Here, the “Ensemble” score was the standard $s(x,y):=\frac{|\mu(x)-y|}{\sigma(x)}$, where $(\mu(x),\sigma(x))$ are the ensemble mean and standard deviation. Prediction intervals could be analytically constructed for the $\mathcal{C}^{M}$, $\mathcal{C}^{R}$, and $\mathcal{C}^{U}$ methods. To assess CSA, however, a discretized grid $\mathcal{G}_{Y}\subset\mathcal{Y}$ of coarseness $\Delta y$ was considered, and an interval length estimate given by $\mathcal{L}(\mathcal{C}(x))\approx\Delta y\cdot|\{y:y\in\mathcal{G}_{Y},s(x,y)\in\widehat{\mathcal{Q}}\}|$. We also present an ablation, labeled “Single-Stage,” to demonstrate the two-stage calibration is necessary to retain coverage; this single-stage approach does not split $\mathcal{S}_{C}$ and instead directly computes $\{\widehat{q}_{m}\}$ on $\mathcal{S}_{C}$ per Section 3.1.1.
We provide the results for $\alpha=0.05$ and $\alpha=0.025$ to demonstrate the consistency of the method performance. A subset of the results is given in Table 2; the full set of results is deferred to Appendix G. As in the results of Section 4.1, we see that CSA retains the coverage guarantees typical of conformal prediction yet produces significantly smaller prediction intervals than both the individual models and the alternate aggregation strategies. We additionally see that the “Single-Stage” approach fails to retain coverage, demonstrating the necessity of the two-stage calibration. We provide a visual comparison of the prediction regions resulting from these methods in Appendix H.
We additionally assessed the robustness of our method to imbalanced ensembles. The experiments of [29] were conducted on a UCI benchmark task [44] with an ensemble of an OLS model, a LASSO linear model, a random forest, and an MLP, and they found the conformalized random forest to outperform all the proposed aggregation strategies, due to the lack of orthogonal information in considering the other predictors. We find that, in these degenerate cases, where the best decision is to simply choose a single predictor, our method outperforms other aggregation methods and nearly matches the performance of the best conformalized predictor in hindset; the results are presented in Appendix I across a number of UCI benchmarks.
Table 2: The results for five distinct tasks are shown below for $\alpha=0.05$ (top five rows) and $\alpha=0.025$ (bottom five rows). For each, the average coverages (grey rows) and prediction set lengths (white rows) with standard deviations are given, both assessed over 5 randomized draws of the training, calibration, and test sets. In cases where the method failed to achieve sufficient coverage (i.e. <.$<.93$ for $\alpha=0.05$ and $<0.96$ for $\alpha=0.025$), we do not include it in comparison for set length.
Dataset/$\boldsymbol{\alpha}$ OLS LASSO RF XGBoost $\boldsymbol{\mathcal{C}^{M}}$ $\boldsymbol{\mathcal{C}^{R}}$ $\boldsymbol{\mathcal{C}^{U}}$ Ensemble Single-Stage CSA
361234 0.97 (0.011) 0.966 (0.011) 0.939 (0.002) 0.954 (0.006) 0.956 (0.011) 0.948 (0.01) 0.96 (0.013) 0.95 (0.006) 0.955 (0.013) 0.957 (0.01)
$(\alpha=0.05)$ 9.673 (0.160) 9.645 (0.154) 10.080 (0.160) 9.157 (0.052) 9.196 (0.123) 8.703 (0.086) 9.524 (0.056) 17.759 (0.275) 7.646 (0.073) 7.688 (0.181)
361235 0.947 (0.0) 0.945 (0.005) 0.968 (0.016) 0.95 (0.005) 0.955 (0.016) 0.897 (0.005) 0.953 (0.011) 0.932 (0.021) 0.745 (0.011) 0.984 (0.005)
$(\alpha=0.05)$ 20.961 (0.651) 24.241 (0.246) 10.096 (0.587) 11.387 (0.452) 11.782 (0.057) 16.088 (0.118) 15.823 (1.272) 6.162 (0.458) 11.695 (0.266)
361236 0.975 (0.008) 0.975 (0.008) 0.961 (0.0) 0.948 (0.012) 0.948 (0.012) 0.938 (0.012) 0.965 (0.008) 0.934 (0.004) 0.94 (0.004) 0.963 (0.004)
[…]
361237 0.969 (0.023) 0.969 (0.023) 0.981 (0.0) 0.923 (0.0) 0.954 (0.015) 0.9 (0.008) 0.969 (0.023) 0.885 (0.038) 0.8 (0.015) 0.977 (0.008)
$(\alpha=0.05)$ 44.019 (0.990) 44.069 (1.115) 27.035 (1.014) 26.524 (1.244) 31.967 (1.118) 14.473 (0.503) 23.145 (0.199)
361241 0.954 (0.001) 0.956 (0.001) 0.944 (0.005) 0.957 (0.002) 0.954 (0.002) 0.923 (0.0) 0.952 (0.0) 0.949 (0.001) 0.917 (0.006) 0.951 (0.001)
$(\alpha=0.05)$ 19.133 (0.062) 20.245 (0.095) 18.102 (0.055) 18.482 (0.062) 17.958 (0.062) 18.932 (0.034) 29.548 (0.191) 15.199 (0.427) 17.328 (0.097)
361234 0.987 (0.008) 0.987 (0.008) 0.974 (0.004) 0.977 (0.008) 0.982 (0.008) 0.971 (0.01) 0.981 (0.01) 0.97 (0.008) 0.976 (0.01) 0.973 (0.006)
$(\alpha=0.025)$ 11.939 (0.137) 11.871 (0.084) 12.484 (0.168) 11.972 (0.009) 11.587 (0.110) 11.157 (0.086) 11.965 (0.050) 25.598 (0.974) 9.306 (0.259) 8.855 (0.059)
361235 0.987 (0.0) 0.982 (0.011) 0.979 (0.011) 0.984 (0.005) 0.989 (0.005) 0.966 (0.016) 0.976 (0.005) 0.958 (0.021) 0.889 (0.011) 0.989 (0.005)
$(\alpha=0.025)$ 24.595 (0.825) 28.841 (1.129) 11.811 (0.992) 14.237 (0.786) 14.472 (0.172) 12.278 (0.026) 19.231 (0.356) 7.719 (0.467) 12.563 (0.766)
361236 0.992 (0.004) 0.992 (0.004) 0.981 (0.0) 0.965 (0.008) 0.975 (0.008) 0.965 (0.008) 0.977 (0.012) 0.955 (0.008) 0.955 (0.012) 0.973 (0.004)
[…]
361237 0.981 (0.0) 0.981 (0.0) 0.981 (0.0) 0.977 (0.008) 0.962 (0.0) 0.962 (0.0) 0.977 (0.008) 0.965 (0.008) 0.927 (0.031) 0.981 (0.0)
$(\alpha=0.025)$ 47.738 (0.542) 47.440 (0.959) 30.785 (0.037) 26.208 (0.897) 30.554 (0.561) 27.182 (0.803) 35.982 (0.619) 67.660 (6.380) 18.214 (0.436) 26.897 (0.515)
361241 0.979 (0.001) 0.978 (0.001) 0.976 (0.001) 0.978 (0.001) 0.978 (0.0) 0.964 (0.002) 0.977 (0.0) 0.972 (0.002) 0.958 (0.003) 0.979 (0.0)
$(\alpha=0.025)$ 21.772 (0.085) 23.089 (0.106) 21.543 (0.009) 21.454 (0.109) 20.862 (0.088) 19.291 (0.060) 21.905 (0.041) 40.082 (0.045) 17.765 (0.329) 19.897 (0.062)
@@ 4.3 CSA Predict-Then-Optimize
We now study a real-world predict-then-optimize traffic routing task, from [36]. In this task, a time series of $T$ preceding precipitations is used to predict future precipitations and, in turn, future traffic, as fully described in Appendix J. We consider the traffic routing problem for a fixed source-target pair $(s,t)$ over the graph of Manhattan, where $|\mathcal{V}|=4584$ and $|\mathcal{E}|=9867$. Formally,
$w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}\widehat{c}^{T}w\quad\textrm{s.t.}\quad w\in[0,1]^{\mathcal{E}},Aw=b,\mathcal{P}_{X,C}(C\in\mathcal{C}(X))\geq 1-\alpha$
@@
We then consider two probabilistic models for traffic prediction, namely one based on the classical probabilistic Lagrangian integro-difference approach (STEPS) of [45] and one on the modern latent diffusion model (LDM) approach of [46]. As a result of the higher inference cost of the latter, we consider the setup where $J_{1}>J_{2}$, specifically with $J_{1}=4$ and $J_{2}=1$, highlighting the flexibility of non-uniform sampling from predictors discussed in Section 3.2. As discussed in Section 4, the alternate aggregation strategies do not lend themselves for use in this setting. We, therefore, only compare CSA to the separate conformalizations of the two predictors, with the score from Equation 2. We here evaluate the methods using the expected suboptimality gap proportion, $\Delta_{\%}=\mathbb{E}_{X}[\Delta(X,C(X))/\min_{w}f(w,C(X))],$ where $\Delta$ is defined as discussed in Section 2.3. This measures the conservatism of the robust optimal value and is bounded in $[0,1]$.
Experiments were conducted with $|\mathcal{D}_{C}|=200$, with a $20/80\%$ split used for $\mathcal{D_{C}}^{(1)}$-$\mathcal{D_{C}}^{(2)}$. The suboptimality was then computed across 100 i.i.d. test samples. To assess the improvement, we conducted two paired t-tests, where $H_{0}:\Delta^{(\mathrm{CSA})}_{\%}=\Delta^{(\mathrm{STEPS})}_{\%}$ and $H_{1}:\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$ and similarly for $\Delta^{(\mathrm{CSA})}_{\%}$ and $\Delta^{(\mathrm{LDM})}_{\%}$. The results are provided in Table 3, from which we find that CSA significantly reduces the suboptimality after accounting for Bonferroni multiple testing. We see that, while conformalization of either of the two views individually already produces the desired coverage, CSA produces more informative prediction regions, and hence less conservative robust upper bounds.
Table 3: Coverages for $\alpha=0.05$ for the individually conformalized and CSA approach and p-values of the paired t-tests comparing $\Delta_{\%}$ are shown, both computed over 100 i.i.d. test samples.
Coverage P-values for $H_{1}$
STEPS LDM CSA $\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$: $3.61\times 10^{-4}$
0.981 0.962 0.968 $\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{LDM})}_{\%}$: $9.50\times 10^{-4}$
@@ 5 Discussion
We have presented a framework for producing informative prediction regions in ensemble predictor pipelines, suggesting many directions for extension. One is in extracting insights of the relative predictor uncertainties from the data-driven $\lesssim$. Another is the integration of CSA with [47], which proposed an end-to-end extension to [36]. Finally, given the prevalence of sensor fusion in robotics, another avenue is to study the use of CSA in robust control.
@@ Appendix C CSA Predict-Then-Optimize Visual Walkthrough
Notice that, for any fixed $\vec{j}$, this is a standard convex optimization problem with a convex feasible region. We illustrate how the feasible region would be constructed for a fixed $\vec{j}$ in Figure 8. Note that the construction of this feasible is never explicitly done in practice and is only implicitly used by convex solver routines in practice. We can, therefore, then solve Equation 10 over all possible $\vec{j}\in\mathcal{J}$ and aggregate the maxima to compute $c^{*}$.
@@
C.1 CSA Predict-Then-Optimize Algorithm
We provide the condensed presentation of the predict-then-optimize algorithm below.
Algorithm 2 Predict-Then-Optimize Under CSA
3:  $w^{(0)}\sim U(\mathcal{W})$
4:  for $t\in\{1,\ldots T\}$ do
5:     for $\vec{j}\in\mathcal{J}$ do $c_{\vec{j}}^{*}\leftarrow\operatorname*{arg\,max}_{c}f(w^{(t)},c)\qquad\mathrm{s.t.}\quad\forall m\in 1,...,M\qquad u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}$
8:  end for
9:  Return $w^{(T)}$
Appendix D Computational Efficiency
D.1 Vectorized Score Computation
We now discuss the vectorized form of the computations discussed in Section 3.1.2. In particular, putting the scores into a matrix $S_{C}^{(1)}=[s_{1},...,s_{N_{C_{1}}}]^{\top}\in\mathbb{R}^{N_{C_{1}}\times K}$ and directions into a matrix $U=[u_{1},...,u_{M}]^{\top}\in\mathbb{R}^{K\times M}$, all the projections $u_{m}^{\top}s_{i}$ can be computed as $S_{C}^{(1)}U\in\mathbb{R}^{N_{C_{1}}\times M}$, where $[S_{C}^{(1)}U]_{i,m}$ is precisely $u_{m}^{\top}s_{i}$. $\tilde{q}\in\mathbb{R}^{M}:=\{\tilde{q}_{m}:=\mathrm{quantile}([S_{C}^{(1)}U]_{:,m};1-\alpha)\}$ is then the quantile per row.
For any test point $s\textquoteright\in\mathbb{R}^{K}$, we can then very efficiently check if it falls into the region by checking if it satisfies $Us\textquoteright\leq\tilde{q}$ component-wise. Each iteration of the loop to find $\beta^{*}$, therefore, is very fast, and we find the search typically converges in 5-10 iterations.
The final step is computing $\widehat{t}$. Computing this follows similarly to above, where we take the scores $S_{C}^{(2)}$, compute projections $S_{C}^{(2)}U\in\mathbb{R}^{N_{C_{2}}\times M}$, find $\widetilde{T}:=S_{C}^{(2)}U/\tilde{q}\in\mathbb{R}^{N_{C_{2}}\times M}$, where division is interpreted as being defined component-wise along the rows, computing the maxima similarly along the rows $[T^{*}]_{i}:=\max\widetilde{T}_{i,:}$, and finally computing $\widehat{t}$ as the $1-\alpha$ quantile of $T^{*}$.
D.2 Empirical Efficiency Validation
To demonstrate the computational efficiency of this vectorized approach, we reran the experiment on the “Parkinsons” UCI task with both varying numbers of predictors ($K$) and projection directions ($M$); for each combination, we measured the total time taken to compute the quantile (i.e. to run Algorithm 1) and to perform the projection to assess coverage for the test points. The additional predictors were taken to be random forests with different numbers of trees. $K$ is given in the left column and $M$ in each column heading, with the entry for each $(K,M)$ pair being reported in seconds. As expected, by the vectorized nature of the computations, as discussed in the main paper, the performance scales gracefully over $M$ at roughly $\mathcal{O}(M)$ and remains roughly constant in $K$.
Table 4: Performance values for varying $K$ (number of predictors) and $M$ (number of projection directions). All values are reported in seconds.
$K$ $M$ 10 100 1000 10000
6 0.111668 0.373029 2.32803 37.2561
8 0.0961056 0.327051 2.16211 36.7216
10 0.117146 0.373464 2.73875 37.2603
12 0.123772 0.384527 2.35386 37.0735
Appendix E Prediction Region Intuition
We now consider a simplified setting of the general procedure to gain insight into the efficiency resulting prediction regions. In particular, we consider a scalar regression setting with $K$ predictors $f_{1},...,f_{K}:\mathcal{X}\rightarrow\mathbb{R}$. We further suppose, with a slight abuse of notation, $f(x)-y\sim\mathcal{N}(0,\Sigma)$ for $f(x):=[f_{1}(x),...,f_{K}(x)]$ and that the scores are taken to be $s_{k}(x,y):=(f_{k}(x)-y)^{2}$. Then, as $\widehat{\mathcal{Q}}$ is precisely the region with minimal volume that captures $1-\alpha$ density, it is precisely given by $\chi_{K,1-\alpha}^{2}$, the $1-\alpha$ quantile of the $\chi^{2}$ distribution with $K$ degrees of freedom. That is, the prediction regions are $\mathcal{C}^{\mathrm{CSA}}(x):=\{y:(f(x)-y)^{\top}\Sigma^{-1}(f(x)-y)\leq\chi_{K,1-\alpha}^{2}\}$. Notably, if $\Sigma=\mathrm{diag}(\{\sigma_{i}^{2}\})$,
$\mathcal{C}^{\mathrm{CSA-diag}}(x):=\left\{y:\sum^{K}_{i=1}\left(\frac{f_{k}(x)-y}{\sigma_{k}}\right)^{2}\leq\chi_{K,1-\alpha}^{2}\right\}.$ (11)
Two insights arise from this expression. The first is that, unlike in the prediction regions for the case of the univariate score function $s_{k}(x,y):=(f_{k}(x)-y)^{2}$, the size $|\mathcal{C}^{\mathrm{CSA-diag}}(x)|$ is dependent on $x$. In the univariate case, the size is $2\widehat{q}_{k}$ across all $x$. Here, however, the feasible set of $y$ becomes smaller the more distinct the values $f_{k}(x)$ are. The second insight, therefore, is that, under such independence of residuals, prediction region sizes are minimized in having well-separated predictions, which suggests that efficiency is optimized by having an ensemble of predictors that learn distinct maps from $\mathcal{X}\to\mathcal{Y}$, such as those that focus on distinct views of the covariate space.
Appendix F Full ImageNet Results
Here we present the complete collection of results for the classification task across additional $\alpha$ than those that could fit in the main paper.